OpenStudy (anonymous):

A student in Greece discovers a pottery bowl that contains 48% of its original amount of C-14. Find the age of the pottery bowl to the nearest year. N=N0e*kt N0= initial amount of C=14 (at time t=0) N= amount of C-14 at time t t= time, in years k=0.0001

OpenStudy (anonymous):

So you know that N/N0 = .48 and k=.00001 (and e=2.718) Your equation becomes: .48 = 2.718*(.00001*t) NOTE: you probably mean \[e ^{kt}\] right?

OpenStudy (anonymous):

Yes i think

OpenStudy (anonymous):

then it would be \[.48=(2.718)^{.00001t}\]

OpenStudy (anonymous):

Does that help? Or do you still need help?

OpenStudy (anonymous):

It says e^-kt actually

OpenStudy (anonymous):

ok, so \[.48=2.718^{-.0001t}\]

OpenStudy (anonymous):

do you know how to solve for t?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

ok, so, I'm guessing that you're in pre-calc or something? to remove e, you have to take the natural log (ln) of both sides of the equation

OpenStudy (anonymous):

I got -7340.4528405

OpenStudy (anonymous):

Let me check

OpenStudy (anonymous):

I got 7339.691 (probably just rounding difference, and don't forget the negative sign on k)

OpenStudy (anonymous):

Wait so it's positive?

OpenStudy (anonymous):

And would that round up to 7340?

OpenStudy (anonymous):

yep, because you can't have negative years, right? So it would be 7340 years. (I think)

OpenStudy (anonymous):

yes, because at 7339 years, it wouldn't have degraded all the way yet. So 7340 is the right answer

OpenStudy (anonymous):

That makes sense, haha. Thank you for helping. I'm awful when it comes to turning a word problem into an equation.

OpenStudy (anonymous):

no problem! glad to help

OpenStudy (anonymous):

would you be willing to help me on 3 others..?

OpenStudy (anonymous):

sure, shoot.

OpenStudy (anonymous):

An archaeologist in Turkey discovers a spear head that contains 63% of its original amount of C-14. Find the age of the spear head to the nearest year.

OpenStudy (anonymous):

With the same equation?

OpenStudy (anonymous):

and the same k?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Well, what would N/N0 be in this equation?

OpenStudy (anonymous):

It's the same formula, I just can't make an equation. If you showed me I could solve it.

OpenStudy (anonymous):

I know, I'm trying to get you to do it. In the last equation, we knew that there was 48% of the original substance remaining, so we knew that Ncurrent/Noriginal (N/N0) was .48 or 48%. So what would N/N0 be in this case where there is 68% of the original substance remaining?

OpenStudy (anonymous):

.68

OpenStudy (anonymous):

right, so just solve for .68 in the same equation \[.68=e ^{-.0001t}\] where e is still the constant 2.718

OpenStudy (anonymous):

understand how I got that?

OpenStudy (anonymous):

Well kinda, I don't understand the 2.718..?

OpenStudy (anonymous):

But I got 3857.02

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/E_%28mathematical_constant%29 e is a mathematical constant, (like pi=3.1415926....) it's based on circle geometry. It's safe to estimate e=2.718

OpenStudy (anonymous):

Oh! Ok thanks I didn't know that.

OpenStudy (anonymous):

No problem, does that help you with the other ones? (I got 3856.625, so 3857, don't forget to round up to the nearest whole number!)

OpenStudy (anonymous):

That does, but hang on the other ones might be different..

OpenStudy (anonymous):

Would 14=2.718^(0.1481t) simplify to 18?

OpenStudy (anonymous):

If you're still rounding up, then yes.

OpenStudy (anonymous):

BUT

OpenStudy (anonymous):

In this case, you have N0=14. What is the problem asking you to solve for?

OpenStudy (anonymous):

The substances half life in days. And to round to the nearest 10th.

OpenStudy (anonymous):

So you want to know how long it takes to get to N=7 (where only half of the 14 grams are left) Can you fix your equation?

OpenStudy (anonymous):

7=2.718^(0.1481t)

OpenStudy (anonymous):

mm, close, you forgot N0

OpenStudy (anonymous):

Ok then I'm confused ;_;

OpenStudy (anonymous):

Ok, so you have N=N0*e^(-kt) right? You know that you start with 14 grams so: N=14*e^(-kt) Then you know that you want to know how long it takes to get to half of 14 grams (=14/2=7) so: 7=14*e^(-kt) Plug in your constants: 7=14*2.718^(-.1481t) Yes?

OpenStudy (anonymous):

Oh, ok. I will solve now.

OpenStudy (anonymous):

4.68?

OpenStudy (anonymous):

yes, which would round to .....? (unless you're allowed to do decimals)

OpenStudy (anonymous):

Well it said on this one to round to the nearest tenth.

OpenStudy (anonymous):

oh right, so 4.7

OpenStudy (anonymous):

.68 would be to the nearest hundredth

OpenStudy (anonymous):

Oh ok :) Thank you.

OpenStudy (anonymous):

no problem, anything else?

OpenStudy (anonymous):

Maybe, not sure yet.

OpenStudy (anonymous):

I will post on here if I have something else.

OpenStudy (anonymous):

Ok, I have to get back to work. I'll try to check back, but no promises. You clearly get it, just keep truckin!